A card is drawn at random from a deck of 52 playing cards. Find the
probability that it is a club or a king.

Found 2 solutions by stanbon, jorel1380:
Answer by stanbon(75887)

A card is drawn at random from a deck of 52 playing cards. Find the
probability that it is a club or a king.
----------------
# of (club OR king) = 13+4-1 = 16
---
P(club or king) = 16/52 = 4/13
-----------------------------------
Cheers,
Stan H.

Answer by jorel1380(3719)

13 clubs + 4 kings - 1 king of clubs=16
kings or clubs =16/52=4/13..